Two point charges are located on the x axis of a coordinate system q1=1.0nC is at x=+2.0cm, and q2=-3.0nC is at x=+4.0cm. what is the total electric force exerted by q1 and q2 on a charge q3=5.0nC at x =0?
1
Expert's answer
2019-11-18T10:44:55-0500

The total electric force on the third charge caused by charges 1 and 2 is the sum of the forces of both charges (1 and 2) acting on the charge 3 located at the beginning of coordinate (i.e. at "x_3=0"):


"F_T=\frac{1}{4\pi\epsilon_0}\frac{q_1q_3}{(x_1-x_3)^2}+\frac{1}{4\pi\epsilon_0}\frac{q_2q_3}{(x_2-x_3)^2}=\\
\space\\
=1.12\cdot10^{-4}\text{ N}."

Substitute charges with their signs and don't forget that nano- means you must multiply by "10^{-9}"!